OpenStudy (anonymous):

z+5 3z-1 4z-4" it's in a shape of a triangle what do i do?

jimthompson5910 (jim_thompson5910):

draw out a pic please

OpenStudy (anonymous):

number 18

jimthompson5910 (jim_thompson5910):

# 18 ??

jimthompson5910 (jim_thompson5910):

oh nvm...lol

OpenStudy (anonymous):

yah lol

jimthompson5910 (jim_thompson5910):

the tick marks mean that each side is congruent to each other

jimthompson5910 (jim_thompson5910):

so z+5 = 4z - 4

jimthompson5910 (jim_thompson5910):

solve for z, then use this to find the length of each side

OpenStudy (anonymous):

so x + 5 = 4z - 4 if you do 4z - 4 u get z. then z +5 = 5z so 5z = z?

jimthompson5910 (jim_thompson5910):

not sure where you got x

OpenStudy (anonymous):

i me z*

OpenStudy (anonymous):

mean*

jimthompson5910 (jim_thompson5910):

but yes, z + 5 = 4z - 4 isolate z to get z + 5 = 4z - 4 5+4 = 4z-z Keep going...

OpenStudy (anonymous):

so 9 = 4z?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

4z - z isn't 4z

OpenStudy (anonymous):

what would it be? im enw to this.

jimthompson5910 (jim_thompson5910):

think of z as 1z

jimthompson5910 (jim_thompson5910):

so 4z - z is the same as 4z-1z

OpenStudy (anonymous):

so wait if youu do 4z - z witch is like 4z - 1z and u subtract that you would get 3z?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so 9 = 3z or 3z = 9

OpenStudy (anonymous):

thats the answer? haha

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

we want z = ... not 3z = ...

OpenStudy (anonymous):

how would you get the from 3z?

jimthompson5910 (jim_thompson5910):

3z means 3 times z

jimthompson5910 (jim_thompson5910):

undo the multiplication of 3 to get ???

OpenStudy (anonymous):

soo sorry if i sound stupid but 3 x z?

jimthompson5910 (jim_thompson5910):

well you divide both sides by 3

OpenStudy (anonymous):

so 3 divided by 3 = 1 and 3 divided by z = 3z?

jimthompson5910 (jim_thompson5910):

3z = 9 z = 9/3 z = 3

OpenStudy (anonymous):

oh!!!!!!!! i get it!!!! so z = 3?

jimthompson5910 (jim_thompson5910):

yes, now use this to find the length of each side

OpenStudy (anonymous):

so eahc side length would be 8 if im correct?

jimthompson5910 (jim_thompson5910):

you are

OpenStudy (anonymous):

thanks :)

jimthompson5910 (jim_thompson5910):

np